OpenStudy (faman39):

A bag contains 3 black marbles and 2 white marbles. Jake reaches in and selects one marble without looking. What is the probability that he gets a white marble? a. 20% b. 40% c.50% d.67%

OpenStudy (ash2326):

@Faman39 Did you try this?

OpenStudy (faman39):

I think, its 40?

OpenStudy (ash2326):

How did you proceed with the problem?

OpenStudy (unklerhaukus):

2/5=?%

OpenStudy (faman39):

0.4

OpenStudy (unklerhaukus):

2/5=4/10=40/100=40%

OpenStudy (unklerhaukus):

right

OpenStudy (ash2326):

Yeah:D